Which of the ordered pairs
(0, 0), (–2, 10), (–1, –5), (–3, 9)
are solutions for the equation y = 5x?

Answer by Edwin McCravy(20056)


Which of the ordered pairs 
(0, 0), (–2, 10), (–1, –5), (–3, 9)
are solutions for the equation y = 5x?


Plug (x,y) = (0,0) into y = 5x

y = 5x
0 = 5(0)
0 = 0

Yes that equation 0 = 0 is true, so (0,0) is a solution.

-------

Plug (x,y) = (-2,10) into y = 5x

y = 5x
10 = 5(-2)
10 = -10

No, that equation 10 = -10 is false, so (0,0) is not a solution.

-------

Plug (x,y) = (-1,-5) into y = 5x

y = 5x
-5 = 5(-1)
-5 = -5

Yes that equation -5 = -5 is true, so (-1,-5) is a solution.

-------

Plug (x,y) = (-3,9) into y = 5x

y = 5x
9 = 5(-3)
9 = -15

No, that equation 9 = -15 is false, so (-3,9) is not a solution.

Edwn